QUESTION 3 Calculate the degree of unsaturation for the following molecular formulas CioH16
Answer #1

Degree of unsaturation:

DOU = (2C + 2 + N - X - H)/2

X is number of halogens

C is number of carbon

N is number of Nitrogen

H is number of Hydrogen

a)

DOU = (2C + 2 + N - X - H)/2

= (2*10 + 2 + 0 - 0 - 16)/2

= (22 - 16) / 2

= 6/2

= 3

Answer: 3

b)

DOU = (2C + 2 + N - X - H)/2

= (2*5 + 2 + 0 - 1 - 5)/2

= (12 - 6) / 2

= 6/2

= 3

Answer: 3

c)

DOU = (2C + 2 + N - X - H)/2

= (2*14 + 2 + 0 - 1 - 29)/2

= (30 - 30) / 2

= 0/2

= 0

Answer: 0

d)

DOU = (2C + 2 + N - X - H)/2

= (2*9 + 2 + 1 - 0 - 15)/2

= (21 - 15) / 2

= 6/2

= 3

Answer: 3